acid in a acne cleanser For each aliquot: 1) Use the calibration curve to determine the concentration of the salicylic acid in the acne cleanser solution through interpolation of the absorbance measured for the acne cleanser sample, by using the equation of the line. 2 Calculate the moles and mass of salicylic acid present in the 25.00 mL volumetric flask. 3Y Determine the amount (mass) of salicylic acid in 1.00 mL of acne cleanser. 4) Using the density from the table below determine the mass of 1.00 mL of acne cleanser Calculate the%mass of salicyclic acid in the cleanser using the 5) equation: (mass salicylic acid/mass acne cleanser) x 100% % mass Determine the average % mass of the salicylic acid in the acne cleanser, the standard deviation of the %mass, and the %RSD. Compare the experimental value of the average%mass with the advertised amount (on the bottle) using the following equation: % comparisons (average % mass of salicylic acid experimentally determined / advertised % mass of salicylic acid) x 100%. Table 2. Density of Acne Cleansers 0,28339 Density (8/ml) Acne Cleanser Clean and Clear Blackhead Clearing Astringent 0.908 0.961 Clean and Clear Deep Cleansing Astringent Iron nitrate solutian Mo.OwM Acne deansen use 0.3ml
Table 3.1. Summary for preparation of solutions in 25.00 mL flasks each containing 5.0 mL of stock iron solution Blank | Std. #1 | Std. #2 | Std. #3 | Unknown (mL) (mL) (mL (mL)(mL) stock salicylic acid 0.0 1 1.00 5.00 10.00 solution 1.00 Acne cleanser o1 017 ,38 .768 .300 Absapbance af 3Snm
Answer #1

Molarity of stock salicylic acid = 0.002036 M i.e, 0.002036 moles/L

Molar mass of salicylic acid = 138.12 g/mol

Answer 1 and 2

As per the table of standard and unknown solution

Solution Volume of salicylic acid added Volume of acne cleanser Moles of Salicylic acid = Molarity * Volume of salicylic acid/1000 Mass of salicylic acid = Moles of salicylic acid * Molar mass of salicylic acid Absorbance at 535 nm
Blank 0.00 mL ----- 0.00 0.00 0.00
Std 1 1.00 mL ----- 0.000002036 mol 0.0002812 g 0.017
Std 2 5.00 mL ------ 0.00001018 mol 0.0014061 g 0.382
Std 3 10.00 mL ------ 0.00002036 mol 0.002812 g 0.768

Standard curve is as below:

0.9 y-271.15x R2 0.991 0.8 0.768 0.7 0.6 0.5 XT 0.4 0.382 0.3 0.2 0.1 0.017 0 0 0.0005 0.001 0.0015 0.002 0.0025 0.003 Mass o

From equation of line, y = 271.15x

Substituting value of absorbance for acne cleaner sample, we get 1.300 = 271.15 x

x = 0.004794 g

Answer 3

0.004794 g is the mass of 0.3 mL acne cleanser (as is handwritten in question)

So, mass of salicylic acid in 1 mL of cleanser = 0.004794/0.3 = 0.01598 g

Answer 4

Mass of acne cleanser from density measurements = Volume * Density of deep cleanser = 1 mL * 0.961 g/mL = 0.961 g

Mass of acne cleanser from density measurements = Volume * Density of blackhead cleanser = 1 mL * 0.908 g/mL = 0.908 g

Answer 5

% mass = (0.01598/0.961)*100 = 1.663 % [For deep cleanser]

% mass = (0.01598/0.908)*100 = 1.759 % [For blackhead cleanser]

Average % mass = 1.711 %

Standard deviation and % RSD can be calculated only if there are more than one unknown samples.

Also, advertised value is not given, so, comparison cannot be done